OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was 3 times his age. How old is Sam now?

OpenStudy (anonymous):

let sam's age=x sam's coach age=2x 10 years before,sam's age=x-10 10 years before,sam's caoch age=3(x-10)=3x-30 2x-10=3x-30 3x-2x=-10+30 x=20

OpenStudy (anonymous):

Let the age of Coach is : x And that of Sam is : y x = 2y And 10 years ago, Coach was : (x - 10) Sam was: (y - 10) So, x- 10 = 3(y - 10) x - 3y = -20 2y - 3y = -20 y = 20 So, Sam is now 20 years old..

OpenStudy (anonymous):

give me my medal

OpenStudy (anonymous):

aye man my computer froze man thanx

OpenStudy (anonymous):

brb finna shower